OpenStudy (anonymous):

the little town art center charges $21 for adults, $14 for senior citizens, and $6 for children, on sunday afternoon, there were 718 tickets sold for a total of $9486 and there were 44 more childrens tickets sold than adults how many childrens tickets were sold

OpenStudy (anonymous):

Make a system of equations. Let a represent adults, s represent seniors, and c represent children.

OpenStudy (anonymous):

The resulting equations are 21a + 14s + 6c = 9486 a + s + c = 718 c = a + 44

OpenStudy (anonymous):

Try solving that system, and if you find a, s, and c you have the answers.